# Primes and absolutely or non-absolutely irreducible elements in atomic domains

### v1

@@ 4.2. Absolutely irreducible elements
Let $p\in\mathcal{O}_{K}$ be an irreducible element with square-free factorization into prime ideals; $p\mathcal{O}_{K}=P_{1}\cdot\ldots\cdot P_{r}$. (Such an element exists because there are unramified primes and $\mathcal{O}_{K}$ is atomic.) Then let $q=\max_{1\leq j\leq r}[\mathcal{O}_{K}:P_{j}]$. W.l.o.g., $[\mathcal{O}_{K}:P_{1}]=q$. Let $r_{1},\ldots,r_{q}$ be a complete system of residues modulo $P_{1}$, containing one (but not more than one) complete system of residues modulo $P_{i}$ for each $i$ and not containing a complete system of residues modulo any other primes. Set
@@
Then $F$ is absolutely irreducible.
To see this, suppose $F^{m}$ factors as $F^{m}=c_{1}g_{1}\ldots c_{s}g_{s}$ with $c_{i}\in K$, $g_{i}$ monic in $K[x]$, and $c_{i}g_{i}\in\operatorname{Int}(\mathcal{O}_{K})$ irreducible for all $i$. Note that $F^{m}$ is image-primitive, and, therefore, so is $c_{i}g_{i}$ for all $i$, meaning $c_{i}=(\operatorname{fd}(g_{i}))^{-1}$.
Let $v_{k}$ be the normalized essential valuation corresponding to $P_{k}$. Then $v_{1}(c_{i})=-v_{1}(\operatorname{fd}(g_{i}))$, and $\sum_{i=1}^{s}v_{1}(c_{i})=v_{1}(p^{-m})=-m$. So, $\sum_{i=1}^{s}v_{1}(\operatorname{fd}(g_{i}))=m$, which can only happen if each $g_{i}$ is a power of $f$; $g_{i}=f^{m_{i}}$ with $\sum_{i=1}^{s}m_{i}=m$.
Now $c_{i}=(\operatorname{fd}(g_{i}))^{-1}=p^{-m_{i}}$, and, therefore, $c_{i}g_{i}=f^{m_{i}}p^{-m_{i}}=F^{m_{i}}$. As $c_{i}g_{i}$ was assumed irreducible, $m_{i}=1$ follows. ∎
@@ 4.3. Non-absolutely irreducible elements
Let $p\in\mathcal{O}_{K}$ be an irreducible element with square-free factorization into prime ideals; $p\mathcal{O}_{K}=P_{1}\cdot\ldots\cdot P_{r}$. (Such an element exists because there are unramified primes and $\mathcal{O}_{K}$ is atomic.) Then let $q=\max_{1\leq j\leq r}[\mathcal{O}_{K}:P_{j}]$. W.l.o.g., $[\mathcal{O}_{K}:P_{1}]=q$. Let $r_{1},\ldots,r_{q}$ be a complete system of residues modulo $P_{1}$, containing one (but not more than one) complete system of residues modulo $P_{i}$ for each $i$ and not containing a complete system of residues modulo any other primes. Set $g(x)=(x-r_{1})^{2}$ and $h(x)=(x-r_{2})\cdots(x-r_{q})$.
@@
Regarding non-absolutely irreducible elements of $\operatorname{Int}(\mathcal{O}_{K})$, we can likewise generalize examples where the $n$-th power of an irreducible element has factorizations of length other than $n$ (for instance, [24, Example 4.4]) from $\operatorname{Int}(\mathbb{Z})$ to $\operatorname{Int}(\mathcal{O}_{K})$ by using [11, Lemma 3.3] as in the above proof.
@@
Furthermore, every irreducible of the form $a(1+xf(x))$ is prime [1]. The irreducible elements of the form $ax$ are not absolutely irreducible, because
@@
    All the valuation rings that appear as localizations of the integral closure of $D$ are independent of $V$.
@@
Let $K$ be any field in which $-1$ is not a square, that is, a formally real field, for instance, $K=\mathbb{Q}$, and let $X$ be an indeterminate over $K$. Define $D$ and $V$ as the localizations
@@
Note that $(x^{2}+1)$ is indeed a prime ideal of $K[X]$ since $-1$ is not a square in $K$. Both, $D$ and $V$, are one-dimensional Noetherian local integral domains and their quotient field is just the function field $K(X)$. In order to see this for $D$, just note that $K[X^{2},X^{3}]$ is a numerical semigroup algebra and hence one-dimensional and Noetherian.
Moreover, $V$ clearly is a discrete rank one valuation ring with prime element $\pi=X^{2}+1$. The integral closure of $D$ is the valuation ring $K[X]_{(X)}$ that is indeed independent of $V$.
Next we show that $X^{2}+1$ is also a prime element of $R$. For this, we can just argue that it generates the prime ideal $P=(X^{2}+1)_{(X^{2}+1)}\cap R$, where $(X^{2}+1)_{(X^{2}+1)}$ denotes the unique maximal ideal of $V$.
Let $g\in P$. As $g$ is, a fortiori, an element of $D$, we can write it in the form $g=\frac{h}{s}$, where $h$ $s\in K[X^{2},X^{3}]$ and $s\notin(X^{2},X^{3})_{K[X^{2},X^{3}]}$. On the other hand, $g$ is also in $(X^{2}+1)_{(X^{2}+1)}$ and we can therefore write it as $g=\frac{a}{b}(X^{2}+1)$ with $a$$b\in K[X]$ and $b\notin(X^{2}+1)_{K[X]}$.
@@
that we can view as an identity in $K[X]$. Since $X^{2}+1$ does not divide $b$ in $K[X]$, it has to divide $h$ and the unique cofactor is $\frac{sa}{b}\in K[X]$. Since $h\in K[X^{2},X^{3}]$ has no linear term, it follows from $h=\frac{sa}{b}(X^{2}+1)$ that the polynomial $\frac{sa}{b}$ also has no linear term, that is, $\frac{sa}{b}\in K[X^{2},X^{3}]$. By choice, $s\in K[X^{2},X^{3}]\setminus(X^{2},X^{3})_{K[X^{2},X^{3}]}$ and hence $\frac{a}{b}=\frac{1}{s}\cdot\frac{sa}{b}\in D$.
As a last step, we argue that the elements $X^{2}$ and $X^{3}$ are still irreducible in $R$. They are both elements of $(X^{2},X^{3})_{K[X^{2},X^{3}]}\cap R$ and, therefore, non-units. In the following, for an irreducible polynomial $p$ of $K[X]$, we denote by $\mathsf{v}_{p}$ the $p$-adic valuation on $K(X)$.
We carry out the argument for $X^{2}$; it is then analogous for $X^{3}$. So, decompose $X^{2}=f\cdot g$ where $f$$g\in R$. We want to show that either $f$ or $g$ is a unit of $R$. The elements of $K[X^{2},X^{3}]_{(X^{2},X^{3})}$ (and therefore those of $R$) either have $X$-adic valuation 00 or $\geq 2$. Since $\mathsf{v}_{X}(f)+\mathsf{v}_{X}(g)=\mathsf{v}_{X}(X^{2})=2$, we can assume without loss of generality that $\mathsf{v}_{X}(f)=2$ and $\mathsf{v}_{X}(g)=0$.
Furthermore, as elements of $K[X]_{(X^{2}+1)}$, the $(X^{2}+1)$-adic valuation of $f$ and $g$ is $\geq 0$. Hence, the equality $0=\mathsf{v}_{X^{2}+1}(X^{2})=\mathsf{v}_{X^{2}+1}(f)+\mathsf{v}_{X^{2}+1}(g)$ implies that $\mathsf{v}_{X^{2}+1}(g)=0$. To conclude, $g$ is an element of $R$ that is in none of the two maximal ideals of $R$ and therefore a unit. This finishes the example.
@@
Lemma 6.1.
@@
Let $R=\mathbb{Z}[\sqrt{d}]$ where $d<0$ is a square-free composite integer not congruent to 1 modulo 4. Firstly, the element $\sqrt{d}=\sqrt{ec}$ is irreducible in $R$, but not absolutely irreducible, because $(\sqrt{d})^{2}=e\cdot c$ is a non-trivial factorization of $(\sqrt{d})^{2}$ (not necessarily into irreducibles).
@@
[…]
@@
  2. (ii)
[…]
It follows by [10, Theorem 3.1] that $2$ is absolutely irreducible in $R$.
Lastly, it follows by [22, Theorem 25] that every odd prime $p$ such that $p\operatorname{\not|}d$ and $d^{\frac{p-1}{2}}\equiv-1\ (\mathrm{mod}\ p)$, is prime in $R$.
In the next subsection, we discuss absolute irreducibility in Krull monoids.
6.1. Absolutely Irreducible Elements in Krull monoids
@@
For the second property, suppose that there exists an irreducible $U\in\mathcal{B}(G_{0})$ that is of the form $U=g^{2}T$ with $T\in\mathcal{F}(G_{0})$, and that there exist $\mathfrak{p}\neq\mathfrak{q}\in\mathfrak{X}(H)$ such that $[\mathfrak{p}]=[\mathfrak{q}]=g$. We may assume $T=[\mathfrak{r}_{1}]\cdots[\mathfrak{r}_{k}]$ with $\mathfrak{r}_{i}\in\mathfrak{X}(H)$. Let $\mathfrak{a}=\mathfrak{r}_{1}\cdots_{v}\mathfrak{r}_{k}$. Then $\mathfrak{p}^{2}\cdot_{v}\mathfrak{a}$ and $\mathfrak{p}\cdot_{v}\mathfrak{q}\cdot_{v}\mathfrak{a}$ are principal ideals, say $aH=\mathfrak{p}^{2}\cdot_{v}\mathfrak{a}$ and $bH=\mathfrak{p}\cdot_{v}\mathfrak{q}\cdot_{v}\mathfrak{a}$ with $a$$b\in H$. Since $\theta$ is a transfer homomorphism, the irreducibility of $U$ implies that of $a$ and $b$. However $a\mid b^{2}$ and so $b$ is not absolutely irreducible, contradicting our assumption.
@@
for some $1\leq i_{1}\leq\cdots\leq i_{m}\leq l$. However, a simple length comparison shows $\{i_{1},\ldots,i_{m}\}=\{1,\ldots,l\}$ and so $aH=bH$. ∎
Given the previous theorem, and the fact that we are often interested in cases when $G_{0}=G$ and $G_{1}=\emptyset$, it is useful to determine when every irreducible of $\mathcal{B}(G)$ is absolutely irreducible. The equivalence of the first two statements of the following lemma are well-known.
Lemma 6.5.

### v2

@@ 4.2. Absolutely irreducible elements
Let $p\in\mathcal{O}_{K}$ be an irreducible element with square-free factorization into prime ideals; $p\mathcal{O}_{K}=P_{1}\cdot\ldots\cdot P_{r}$. (Such an element exists because there are unramified primes and $\mathcal{O}_{K}$ is atomic.) Then let $q=\max_{1\leq j\leq r}[\mathcal{O}_{K}:P_{j}]$. W.l.o.g., $\mathcal{O}_{K}:P_{1}]=q$. Let $r_{1},\ldots,r_{q}$ be a complete system of residues modulo $P_{1}$, containing one (but not more than one) complete system of residues modulo $P_{i}$ for each $i$ and not containing a complete system of residues modulo any other primes. Set
@@
We will show that $F$ is absolutely irreducible.
Suppose $F^{m}$ factors as $F^{m}=G_{1}\ldots G_{s}$, with each $G_{i}$ irreducible in $\operatorname{Int}(\mathcal{O}_{K})$, and $G_{i}=c_{i}g_{i}$ with $c_{i}\in K$, $g_{i}$ monic in $K[x]$. Since $F^{m}$ is image-primitive, so is $G_{i}$ for each $i$. This means that $c_{i}=(\operatorname{fd}(g_{i}))^{-1}$.
Let $v$ be the essential valuation corresponding to $P_{1}$, normalized to have value group $\mathbb{Z}$. Then, in particular, $v(c_{i})=-v(\operatorname{fd}(g_{i}))$, therefore,
$\sum_{i=1}^{s}v(c_{i})=v(p^{-m})=-m\quad\text{\ and\ }\quad\sum_{i=1}^{s}v(\operatorname{fd}(g_{i}))=m.$
This can only happen if each $g_{i}$ is a power of $f$; $g_{i}=f^{m_{i}}$ with $\sum_{i=1}^{s}m_{i}=m$.
To see this, consider that each $g_{i}$ is a product of monic linear polynomials $(x-r_{j})$, $1\leq j\leq q$, where $r_{1},\ldots,r_{q}$ form a complete system of residues modulo $P_{1}$. For such a polynomial $g=\prod_{j=1}^{q}(x-r_{j})^{k_{j}}$, clearly $v(\operatorname{fd}(g))=\min_{1\leq j\leq q}k_{j}$.
Returning to $F^{m}=c_{1}g_{1}\ldots c_{s}g_{s}$, where $g_{i}=\prod_{j=1}^{q}(x-r_{j})^{k_{ij}}$ then, of course, $\sum_{i=1}^{s}k_{ij}=m$ for each $j$, while, on the other hand,
$m=\sum_{i=1}^{s}v(\operatorname{fd}(g_{i}))=\sum_{i=1}^{s}\min_{1\leq j\leq q}k_{ij}.$
This implies that for each $1\leq i\leq s$ and $1\leq h\leq q$, necessarily $k_{ih}=\min_{1\leq j\leq q}k_{ij}$, so that each linear factor occurs in $g_{i}$ to the same exponent. If $m_{i}=\min_{1\leq j\leq q}k_{ij}$ then $g_{i}=f^{m_{i}}$.
Now $c_{i}=(\operatorname{fd}(g_{i}))^{-1}=p^{-m_{i}}$, and, therefore, $G_{i}=c_{i}g_{i}=f^{m_{i}}p^{-m_{i}}=F^{m_{i}}$. As $c_{i}g_{i}=G_{i}$ was assumed irreducible, $m_{i}=1$ follows. ∎
@@ 4.3. Non-absolutely irreducible elements
Let $p\in\mathcal{O}_{K}$ be an irreducible element with square-free factorization into prime ideals; $p\mathcal{O}_{K}=P_{1}\cdot\ldots\cdot P_{r}$. (Such an element exists because there are unramified primes and $\mathcal{O}_{K}$ is atomic.) Then let $q=\max_{1\leq j\leq r}[\mathcal{O}_{K}:P_{j}]$. W.l.o.g., $\mathcal{O}_{K}:P_{1}]=q$. Let $r_{1},\ldots,r_{q}$ be a complete system of residues modulo $P_{1}$, containing one (but not more than one) complete system of residues modulo $P_{i}$ for each $i$ and not containing a complete system of residues modulo any other primes. Set $g(x)=(x-r_{1})^{2}$ and $h(x)=(x-r_{2})\cdots(x-r_{q})$.
@@
Regarding non-absolutely irreducible elements of $\operatorname{Int}(\mathcal{O}_{K})$, we can likewise generalize examples where the $n$-th power of an irreducible element has factorizations of length other than $n$ (for instance, [24, Example 4.4]) from $\operatorname{Int}(\mathbb{Z})$ to $\operatorname{Int}(\mathcal{O}_{K})$ by using [11, Lemma 3.3] as in the above proof. To summarize:
Theorem 4.9.
For any number field $K$, the ring of integer-valued polynomials on algebraic integers,
$\operatorname{Int}(\mathcal{O}_{K})=\{f\in K[x]\mid f(\mathcal{O}_{K})\subseteq\mathcal{O}_{K}\},$
is a ring without prime elements containing both absolutely irreducible and non-absolutely irreducible elements.
Proof.
The non-existence of primes is shown in Lemma 4.6; the existence of absloutely irreducible elements in Lemma 4.7, and the existence of non-absolutely irreducible elements in Lemma 4.8. ∎
@@
Furthermore, every irreducible of the form $a(1+xf(x))$ is prime [1]. The irreducible elements of the form axax are not absolutely irreducible, because
@@
    All the valuation rings that appear as localizations of the integral closure of $D$ are distinct from $V$.
@@
Let $K$ be any field in which 1-1 is not a square, for instance $K=\mathbb{Q}$, and let $X$ be an indeterminate over $K$. Define $D$ and $V$ as the localizations
@@
Note that $(X^{2}+1)$ is indeed a prime ideal of $K[X]$ since 1-1 is not a square in $K$. Both, $D$ and $V$, are one-dimensional Noetherian local integral domains and their quotient field is just the function field $K(X)$. In order to see this for $D$, just note that $K[X^{2},X^{3}]$ is a numerical semigroup algebra and hence one-dimensional and Noetherian.
Moreover, $V$ clearly is a discrete rank one valuation ring with prime element $\pi=X^{2}+1$. The integral closure of $D$ is the valuation ring $K[X]_{(X)}$ that is indeed distinct from $V$.
Next we show that $X^{2}+1$ is also a prime element of $R$. For this, we can just argue that it generates the prime ideal $P=\mathfrak{m}_{V}\cap R$, where $\mathfrak{m}_{V}$ denotes the unique maximal ideal of $V$.
Let $g\in P$. As $g$ is, a fortiori, an element of $D$, we can write it in the form $g=\frac{h}{s}$, where $h,s\in K[X^{2},X^{3}]$ and $s\notin(X^{2},X^{3})K[X^{2},X^{3}]$. On the other hand, $g$ is also in $\mathfrak{m}_{V}$ and we can therefore write it as $g=\frac{a}{b}(X^{2}+1)$ with $a,b\in K[X]$ and $b\notin(X^{2}+1)K[X]$.
@@
that we can view as an identity in $K[X]$. Since $X^{2}+1$ does not divide $b$ in $K[X]$, it has to divide $h$ and the unique cofactor is $\frac{sa}{b}\in K[X]$. Since $h\in K[X^{2},X^{3}]$ has no linear term, it follows from $h=\frac{sa}{b}(X^{2}+1)$ that the polynomial $\frac{sa}{b}$ also has no linear term, that is, $\frac{sa}{b}\in K[X^{2},X^{3}]$. By choice, $s\in K[X^{2},X^{3}]\setminus(X^{2},X^{3})K[X^{2},X^{3}]$ and hence $\frac{a}{b}=\frac{1}{s}\cdot\frac{sa}{b}\in D$.
As a last step, we argue that the elements $X^{2}$ and $X^{3}$ are still irreducible in $R$. They are both elements of $R\cap(X^{2},X^{3})K[X^{2},X^{3}]$ and, therefore, non-units. In the following, for an irreducible polynomial $p$ of $K[X]$, we denote by $\mathsf{v}_{p}$ the $p$-adic valuation on $K(X)$.
[…]
Furthermore, as elements of $V=K[X]_{(X^{2}+1)}$, the $(X^{2}+1)$-adic valuation of $f$ and $g$ is $\geq 0$. Hence, the equality $0=\mathsf{v}_{X^{2}+1}(X^{2})=\mathsf{v}_{X^{2}+1}(f)+\mathsf{v}_{X^{2}+1}(g)$ implies that $\mathsf{v}_{X^{2}+1}(g)=0$. To conclude, $g$ is an element of $R$ that is in neither of the two maximal ideals of $R$ and therefore a unit. This finishes the example.
@@
Example 6.1.
Let $R=\operatorname{Int}(\mathbb{Z})\subseteq\mathbb{Q}[x]$ be the ring of integer-valued polynomials as in section 4 (or any atomic domain satisfying ACCP and having both absolutely irreducible and non-absolutely irreducible elements).
We show that $T=R[y]$, the polynomial ring in one indeterminate over $R$, is an atomic domain that has primes, as well as both absolutely and non-absolutely irreducible elements.
Since $R$ is an integral domain, $y$ is a prime element of $T$.
Again because $R$ is an integral domain, by degree considerations, the set of polynomials of degree zero in $R[y]$ (which we may identify with $R\setminus\{0\}$) forms a saturated (i.e., divisor closed) multiplicatively closed subset of $R[y]$ containing all the units of $R[y]$.
A constant in $R[y]$, therefore, is a unit, or absolutely irreducible, or non-absolutely irreducible, in $R[y]$ if and only if it has the respective property in $R$.
Since $R$ contains both absolutely and non-absolutely irreducible elements, $R[y]$ contains both absolutely and non-absolutely irreducible elements (namely, the constants with the respective property in $R$), as well as a prime element, $y$.
Also, since $R$ satisfies ACCP, so does $R[y]$, and $R[y]$ is, therefore, atomic.
Proposition 6.2.
@@
Let d=ec,d=-ec, where $e,c\in\mathbb{Z},e,c\geq 2,d$ is square-free and $d\not\equiv 1\ (\mathrm{mod}\ 4).$ Set $R=\mathbb{Z}[\sqrt{d}]$. Firstly, the element $\sqrt{d}=\sqrt{-ec}$ is irreducible in $R$, but not absolutely irreducible, because $(\sqrt{d})^{2}=-e\cdot c$ is a non-trivial factorization of $(\sqrt{d})^{2}$ (not necessarily into irreducibles).
@@
    if $d\equiv 3\ (\mathrm{mod}\ 4)$, then $2\operatorname{|}(1+\sqrt{d})(1-\sqrt{d})$, but 2 does not divide the individual factors.
@@
    $2R=(2,\sqrt{d})^{2}$ if $d\equiv 2\ (\mathrm{mod}\ 4)$, and
  2. (ii)
    $2R=(2,1+\sqrt{d})^{2}$ if $d\equiv 3\ (\mathrm{mod}\ 4)$.
It follows by [10, Theorem 3.1] that $2$ is absolutely irreducible in $R$.
@@
7. Absolutely Irreducible Elements in Krull monoids
The examples in Sections 8 and 9 will be constructed using Dedekind domains. The multiplicative monoid of non-zero elements of a Dedekind domain is a Krull monoid [15, Chapter 2.10]. To lay the groundwork for Sections 8 and 9, we therefore first prove some results on absolute irreducible elements in Krull monoids, culminating in Theorem 7.2, which forms the basis of Propositions 8.1 and 9.1.
As corollaries we also recover a generalization of a theorem of Chapman and Krause (Corollary 7.4) and a theorem of Angermüller (Corollary 7.5). Corollary 7.5 in particular rules out the existence of Krull monoids having non-prime irreducibles, but having every absolutely irreducible prime. This shows that no example as in Section 5 is possible among Krull monoids. Corollary 7.4 shows that, if every class of the class group of a Krull monoid $H$ contains a prime divisor and there exist no non-absolutely irreducible irreducibles, then $H$ is already factorial. This explains why in Sections 8 and 9 we have to construct domains in which the set of classes containing prime divisors is a proper subset of the entire class group.
@@ 7. Absolutely Irreducible Elements in Krull monoids
[…]
@@
Suppose that $b\in H$ is an irreducible with $b\mid a^{n}$ for some $n\geq 1$. Then $\theta(b)\mid\theta(a)^{n}$. Since $\theta(a)$ is absolutely irreducible, we get $\theta(a)=\theta(b)$. Since $\mathfrak{p}_{i}]\in G_{1}$ for all $1\leq i\leq k$, the image $\theta(b)$ fully determines the multiplicity of each $\mathfrak{p}_{i}$ in $bH$. Now
@@
for some $1\leq i_{1}\leq\cdots\leq i_{m}\leq l$. Since $\mathfrak{q}_{i_{r}}]\in G_{0}\setminus G_{1}$ for $1\leq r\leq m$, assumption (ii) applied to $\theta(b)$ implies $\mathfrak{q}_{i_{r}}]\neq[\mathfrak{q}_{i_{s}}]$ for $r\neq s$. Thus $1\leq i_{1}<\cdots<i_{m}\leq l$. Now $|\theta(a)|=|\theta(b)|$ shows $\{i_{1},\ldots,i_{m}\}=\{1,\ldots,l\}$, and so $aH=bH$. ∎
7.1. Consequences for Krull monoids
As stated at the beginning of the section, we now note some easy consequences of Theorem 7.2 that connect to the existing literature and are of interest in their own right. Aside from the fact that Corollary 7.4 can be used to obtain (ii) of Fact 6.3, these results are however not used in this paper.
Given Theorem 7.2, and the fact that we are often interested in cases when $G_{0}=G$ and $G_{1}=\emptyset$, it is useful to determine when every irreducible of $\mathcal{B}(G)$ is absolutely irreducible. The equivalence of the first two statements of the following corollary is well-known.
Corollary 7.3.